Suppose that antitrust authorities are evaluating whether or not to allow a horizontal merger. Before the merger, the market was competitive, and the merging firms supplied a total of 5,000 goods to the market at a price of $40 per good. Of course, this means that the firms average costs were $40 per good pre-merger. After the merger, expert economists predict that the merged firm will be able to cut its average costs down to$38 per good, but their increased market power will also allow them to sell 4,000 goodsat $45 per good.

(a) Describe the Williamson Tradeoff or depict graphically

(b) Calculate the Williamson Tradeoff that results from this merger

(c) Should this merger be allowed?
Answer #1

1. Williamson Trade Off model in economics describes the trade off between the gains resulting from lower cost of production due to synergy and the losses which are associated with the higher prices due to larger degree of monopoly power that occurs when a horizontal merger happens. Horizontal merger happens when two firms within the industry merge into 1.

When the horizontal merger happens, two effects take place

a. reduction in the competition within the industry as the competitors have joined hands. This in turn helps the firm to raise prices as there are no competitions now and the firms have the power to charge as much price as they want

b. This also reduces the cost of production and other operations as this creates synergy.

In this question, the initial price charger was $40 which is depicted by point c in the diagram. After the merger, the marginal costs will fall because of synergy and the price will rise because of monopoly power. The new price of $45 is depicted in the diagram by p'. Aggregate social surplus will be the area ABCA. Social surplus is the sum of consumer surplus and the producer surplus. Aggregate social surplus after the merger will be ADEFA. Now whether the social surplus has increase or decreased depends on the area of the orange triangle and the area of the red rectangle.

2. Area of Triangle is 1/2*base*height = 1/2*1000*5=2500

Area of rectangle is length*breadth= 4000*2=8000

Hence, decrease in the social surplus is 8000-2500= 5500. This is the Williamson Trade Off.

3. Now, the merger should be allowed in 2 situations.

a. the consumer surplus is not reduced.

b. it does not decrease the total social surplus.

Since the social surplus is getting decreased, the merger should not be allowed.